# {dedication}

### v1

@@
Connecting dots faster, one accelerator at a time.
@@
The advent of Graph Neural Networks (GNNs) has revolutionized the field of machine learning, offering a novel paradigm for learning on graph-structured data. Unlike traditional neural networks, GNNs are capable of capturing complex relationships and dependencies inherent in graph data, making them particularly suited for a wide range of applications including social network analysis, molecular chemistry, and network security. The impact of GNNs in these domains is profound, enabling more accurate models and predictions, and thereby contributing significantly to advancements in these fields.
@@
Synthesizing these insights and optimizations, we design a state-of-the-art hardware accelerator capable of efficiently handling various GNN workloads. Our accelerator architecture is built on our characterization of GNN computational demands, providing clear motivation for our approach. Furthermore, we extend our exploration to emerging GNN workloads in the domain of graph neural networks. This exploration into novel models underlines our comprehensive approach, as we strive to enable accelerators that are not just performant, but also versatile, able to adapt to the evolving landscape of graph computing.
@@ 1 Background and Motivation
    Social Network Analysis: Graphs are essential in Social Network Analysis (SNA) to understand interactions within networks such as Facebook and Twitter. By examining these structures, SNA can detect key influencers, community structures, and predict potential trends or misinformation spread [28, 144, 158, 77].
@@
    Computer Networks: Graphs represent devices and communication pathways in computer networks, aiding in tasks such as routing and fault detection. Graph-based algorithms optimize network design and manage potential vulnerabilities [70, 58, 175, 134].
@@
    Hardware Security: Circuits can be represented as graphs to detect anomalies, such as potential Hardware Trojans. Graph Neural Networks (GNNs) further enhance detection capabilities, pinpointing subtle irregularities [4, 65, 165, 125, 97, 126].
@@
    Transportation and Logistics: Graphs map transportation networks, aiding in solving optimization problems such as route planning and vehicle routing. They help ensure efficient deliveries, optimize public transit, and adapt to disruptions [59, 89, 39, 162, 75].
@@
    Physics: Graphs in physics are used to model quantum states, lattice structures, and network systems. They aid in understanding quantum computing, material properties, and system dynamics [27, 154, 12, 68, 128].
@@
    Machine Learning and Data Mining: In machine learning, graphs facilitate clustering, classification, and feature extraction. They model non-Euclidean data, aiding in semi-supervised learning and geometric deep learning [176, 30, 127, 21, 153, 25, 125].
@@
    Telecommunications: Graphs underpin the telecommunications sector, assisting in routing, optimization, and network slicing for 5G technologies [54, 149, 3, 113, 141].
@@ 3 Objectives and Contributions
Analyze the architectural impact of computing Graph Neural Networks.
Provide a benchmark suite for comprehensive evaluation of how different architectural components influence the performance and efficiency of GNN computations.
Accelerate SpGEMM kernel on custom accelerator.
Propose and implement optimization strategies for the SpGEMM kernel, resulting in significant performance improvements on a custom hardware accelerator.
Design a new accelerator to accelerate multiple GNN workloads.
Develop a novel hardware accelerator architecture tailored for various GNN workloads, demonstrating versatility and improved performance across different graph inputs.
@@
4 Thesis Proposal Organization
  • Chapter Enabling Accelerators for Graph Computing: This chapter sets the stage for the thesis by highlighting the significance and impact of computations based on graph structures.
@@ 4 Thesis Proposal Organization
  • Chapter 4: Building on the insights gained from the previous chapters, we propose a new custom hardware solution designed to accelerate GNN workloads. We present details of the design requirements, resulting architecture, and implementation of the solution, describing how we address the unique challenges posed by GNN computations.
  • Chapter 5: The final chapter of the thesis summarizes the key findings and contributions, discussing both the completed aspects of the research and outlining avenues for future work to be completed as a part of the thesis.
Chapter 1 Preliminaries
Before delving into the details of our accelerator proposal for graph computing, we explore the foundational concepts of graph computing and accelerator design. This chapter serves as a compass, offering a systematic overview of the essential background, terminology and challenges that characterize graph computing, as well as its interaction with hardware accelerators. By first reviewing these foundational elements, this chapter aims to provide the necessary context and knowledge required to understand GNN accelerator design.
@@ 3 Graph Neural Network Frameworks
Programming Language
Primarily Python
Primarily Python
Deep Learning Framework
Built on top of PyTorch
Supports both PyTorch and TensorFlow
Ease of Use
High-level API, easy to use for beginners
Low-level C++ API available
Performance
Efficient and scalable
Highly optimized for sparse operations
Models Available
Extensive collection of pre-implemented GNN models
Newer models only optimized for PyG
Community and Support
Large community, active development
Relatively new library with community beginning to grow
Documentation
Extensive documentation with examples
Comprehensive documentation
Graph Types Supported
Supports heterogeneous and temporal graphs
Supports heterogeneous graphs
Scalability
Optimized for single-machine, multi-GPU setups
Designed for distributed training on multiple machines and large graphs
Extensibility
Easy to extend and contribute
Low-level C++ API relatively difficult to extend
@@ 2 Prior GNN Accelerators
Accelerator
Kernel Fusion
Loop Reordering
Pruning
Bank Mapping
Load Balancing
Hash based
@@
ReFlip [73]: ReFlip is a GCN accelerator that aims to improve the overall efficiency of both regular neural network computations and irregular graph analytics. ReFlip employs a unified architecture based on Processing-in-Memory (PIM) and features a crossbar structure. This unified architecture is augmented by novel algorithm mappings that maximize performance by leveraging the inherent parallelism of crossbar structures.
@@
In this thesis proposal , we address this need by presenting GNNMark [13], a feature-rich benchmark suite that encompasses the diversity present in GNN training workloads, datasets and GNN frameworks. Our benchmark suite consists of GNN workloads that utilize various graph-based data structures, including homogeneous graphs, dynamic graphs, and heterogeneous graphs commonly used in a number of application domains that we mentioned in Section 2. We use this benchmark suite to explore and characterize GNN training behavior on GPUs. We study a variety of aspects of GNN execution, including both compute and memory behavior, highlighting major bottlenecks observed during GNN training. At the system level, we evaluate multiple metric, including the scalability of training GNNs across a multi-GPU system, as well as the sparsity of data encountered during training. The insights derived from our work can be leveraged by both hardware and software developers to improve both the hardware and software performance of GNN training on GPUs. The contributions of this part of the thesis include:
@@ 1 Motivation for Characterizing GNN Workloads
Deep Neural Networks (DNNs) have revolutionized numerous areas, such as image classification [44], speech recognition [114] and autonomous systems [102]. Notable DNN architectures such as Convolutional Neural Networks (CNNs) [52] and Transformers [146] primarily operate on Euclidean data. This type of data, inherently 1D or 2D, includes images and speech datasets. Yet, much of the data we encounter in the real world is non-Euclidean in nature [21], encompassing structures of molecules, social networks, sensor systems, and manifolds. Traditional DNNs, designed for Euclidean data, often fall short in efficiently processing non-Euclidean data due to challenges in directly applying operations, such as convolutions [21].
@@
With GPUs establishing themselves as the go-to platform for DNN and GNN training due to their advanced capabilities, many leading GNN frameworks, for example, the Deep Graph Library [173] and PyTorch Geometric [49], have integrated GPU training support. As GNNs continue to surge in popularity, there’s a pressing need for optimizing GPU platforms to train them. Thoroughly analyzing GPU behavior during GNN training is paramount. By dissecting the myriad of GNN operations and their execution during training, GPU architects can pinpoint and address performance bottlenecks. Aspects such as GNN training scalability on multi-GPU setups and the presence of data sparsity during training can be tapped into for training large graphs, especially those surpassing a single GPU’s memory capacity [122]. A thorough analysis of GNN training workflows will enhance our understanding of the computational and memory constraints associated with running GNN workloads on GPUs.

### v2

@@
To my family,
for their endless love, support, and encouragement.
@@
Acknowledgements.
I wish to extend my profound appreciation to my advisor, Dr. David Kaeli, for his pivotal role in my academic journey. His remarkable foresight and unwavering dedication have played a crucial role in shaping my path towards academic excellence. Dr. Kaeli’s profound expertise and innovative spirit have served as a beacon of guidance through challenging periods, illuminating the path with his profound knowledge and strategic insight. His exceptional commitment to research and education has not only inspired me but also significantly influenced my professional development. The supportive and enriching environment he fostered has been instrumental in navigating the complexities of doctoral research, especially as an international student adapting to a new academic and cultural setting. Dr. Kaeli’s mentorship extends beyond academic achievements, providing invaluable lessons in resilience and integrity. I am eternally grateful for the privilege of being mentored by a scientist of such high calibre and integrity, whose influence will undoubtedly echo throughout my career. I extend my heartfelt gratitude to Prof. Ajay Joshi for his invaluable insights on Homomorphic Encryption, Prof. John Kim for his expertise on network topologies, Prof. José Luis Abellán for his deep knowledge on Graph Neural Networks, and Prof. Devesh Tiwari for his guidance on compute micro-architectures. Their contributions have been instrumental in shaping my research and enhancing my understanding of these complex areas. Additionally, my sincere thanks are due to all faculty members and administrative staff of the College of Engineering at Northeastern University for their unwavering support throughout my doctoral journey. My sincere and warm thanks go also to my lab colleagues: Amir Taherin, Derek Rodriguez, Julian Gutierrez, Malith Jayaweera, Matin Raayai, Michael Shen, Neal Livesay, Nicolas Agostini, Sana Anvari, Trinayan Baruah, Yifan Sun, Alexander Ingare, and Zlatan Feric, for their invaluable contribution to my academic and personal growth. Their collaboration, insight, and camaraderie have been instrumental in fostering a nurturing and productive research environment. It has been a privilege to work alongside such dedicated and talented individuals. Their diverse perspectives and expertise have greatly enriched my research experience. Furthermore, their support and encouragement have been a constant source of motivation, making my journey through the intricacies of academic research both rewarding and memorable. A Ph.D. journey that lasted 6 years, 7 months, and 25 days (or 2429 days in total) was made possible by the unwavering support from my friends Ahan Kak and Ashwin Shirsat. Sharing the joys of highs and support in the times of lows. Finally, last but by no means least, words will fail to express my gratitude to my family, who have been my rock through this journey, especially my parents Chandrakant and Anagha, and my brother Saumil Shivdikar. This achievement would have not been possible without their love, support, and encouragement.
The advent of Graph Neural Networks (GNNs) has revolutionized the field of machine learning, offering a novel paradigm for learning on graph-structured data. Unlike traditional neural networks, GNNs are capable of capturing complex relationships and dependencies inherent in graph data, making them particularly suited for a wide range of applications including social network analysis, molecular chemistry, and network security. The impact of GNNs in these domains is profound, enabling more accurate models and predictions, and thereby contributing significantly to advances in these fields.
@@ Acknowledgements.
Synthesizing these insights and optimizations, we design state-of-the-art hardware accelerators capable of efficiently handling various GNN workloads. Our accelerator architectures are built on our characterization of GNN computational demands, providing clear motivation for our approaches. Furthermore, we extend our exploration to emerging GNN workloads in the domain of graph neural networks. This exploration into novel models underlines our comprehensive approach, as we strive to enable accelerators that are not just performant, but also versatile, able to adapt to the evolving landscape of graph computing.
@@ 1 Background and Motivation
    Social Network Analysis: Graphs are essential in Social Network Analysis (SNA) to understand interactions within networks such as Facebook and Twitter. By examining these structures, SNA can detect key influencers, community structures, and predict potential trends or misinformation spread [31, 182, 198].
@@
    Computer Networks: Graphs represent devices and communication pathways in computer networks, aiding in tasks such as routing and fault detection. Graph-based algorithms optimize network design and manage potential vulnerabilities [81, 68, 218].
@@
    Hardware Security: Circuits can be represented as graphs to detect anomalies, such as potential Hardware Trojans. Graph Neural Networks (GNNs) further enhance detection capabilities, pinpointing subtle irregularities [4, 76, 205].
@@
    Transportation and Logistics: Graphs map transportation networks, aiding in solving optimization problems such as route planning and vehicle routing. They help ensure efficient deliveries, optimize public transit, and adapt to disruptions [69, 108, 47, 202].
@@
    Physics: Graphs in physics are used to model quantum states, lattice structures, and network systems. They aid in understanding quantum computing, material properties, and system dynamics [29, 193, 13, 79].
@@
    Machine Learning and Data Mining: In machine learning, graphs facilitate clustering, classification, and feature extraction. They model non-Euclidean data, aiding in semi-supervised learning and geometric deep learning [219, 34, 23, 192].
@@
    Telecommunications: Graphs underpin the telecommunications sector, assisting in routing, optimization, and network slicing for 5G technologies [64, 188, 3, 141].
@@ 3 Objectives and Contributions
Analyze the architectural impact of computing Graph Neural Networks. Provide a benchmark suite for comprehensive evaluation of how different architectural components influence the performance and efficiency of GNN computations.
Accelerate SpGEMM kernel on custom accelerator. Propose and implement optimization strategies for the SpGEMM kernel, resulting in significant performance improvements on a custom hardware accelerator.
Design a new accelerator to accelerate multiple GNN workloads. Develop a novel hardware accelerator architecture tailored for various GNN workloads, demonstrating versatility and improved performance across sparse and dense compute kernels.
@@
4 Dissertation Organization
  • Chapter : This chapter sets the stage for the thesis by highlighting the significance and impact of computations based on graph structures.
@@ 4 Dissertation Organization
  • Chapter 4: We explore the SMASH (Sparse Matrix Atomic Scratchpad Hashing) algorithm, a novel SpGEMM kernel optimization aimed at enhancing GNN processing. We discuss the development and implementation of SMASH, including its various versions tailored to exploit distinct architectural features for improved efficiency in GNN workloads.
  • Chapter 5: In this chapter, we introduce NeuraChip, a custom CGRA-based accelerator designed to meet the unique demands of GNN computations. We provide detailed discussion on the architecture, including its heterogeneous processing approach, adaptive hash-based compute mapping, and mechanisms for rolling evictions, highlighting how NeuraChip addresses critical bottlenecks in GNN acceleration.
  • Chapter 6: The concluding chapter discusses the key insights and contributions of this work, ranging from the development of a GNN benchmark suite, to the introduction of algorithmic and hardware innovations such as SMASH and NeuraChip. We reflect on the impact of these contributions on the field of GNN acceleration. We also cover potential directions for future GNN research to explore further advancements in graph computing.
Chapter 1 Fundamentals of Graph Computing and Accelerator Architectures
Before delving into the details of our accelerator proposal for graph computing, we explore the foundational concepts of graph computing and accelerator design. This chapter serves as a foundation, offering a systematic overview of the essential background, terminology and challenges that characterize graph computing, as well as its impact on hardware accelerators. By first reviewing these foundational elements, this chapter aims to provide the necessary background required to understand GNN accelerator design.
@@ 3 Graph Isomorphism Networks
4 Graph Attention Networks
Graph Attention Networks (GATs) mark a significant evolution in graph neural network technology by introducing an attention mechanism that allows nodes to dynamically assign importance to their neighbors’ information. This attention-based approach enables the model to focus more on relevant features from a neighborhood, enhancing the adaptability and performance of the network on graph-structured data. GATs address the limitation of conventional graph neural networks, such as GCNs, which treat all neighbors equally during the aggregation process. By incorporating attention, GATs can assign a weight based on the influence of each neighbor based on the task at hand, leading to more effective learning outcomes.
The propagation rule for a GAT layer can be described as:
$h^{(l+1)}_{i}=\sigma\left(\sum_{j\in\mathcal{N}(i)\cup\{i\}}\alpha_{ij}^{(l)}W^{(l)}h^{(l)}_{j}\right)$ (4)
Where:
$h^{(l+1)}_{i}$ is the feature vector of node $i$ at layer $l+1$.
$\alpha_{ij}^{(l)}$ represents the attention coefficient between nodes $i$ and $j$ at layer $l$, indicating the significance of node $j$’s features to the update of node $i$’s features.
$W^{(l)}$ is the weight matrix for layer $l$.
$\sigma$ is an activation function, such as the ReLU function.
The computation of attention coefficients involves a self-attention mechanism where a shared attention function, applicable to all edges, computes the coefficients based on the features of the nodes at either end of the edge. This process allows GATs to perform feature extraction that is both context-aware and adaptive, leading to more nuanced representations of nodes based on their local graph topology.
GATs have demonstrated superior performance on various tasks, including node classification, graph classification, and link prediction, particularly in scenarios where the relevance of neighboring nodes varies significantly. The model’s ability to selectively prioritize information makes it highly effective in capturing the complex dependencies characteristic of graph-structured data, thereby pushing the boundaries of what is achievable with graph neural networks.
5 GraphSAGE
GraphSAGE (Graph Sample and AggregatE) is a novel framework designed to efficiently generate node embeddings for large-scale graphs. Unlike traditional graph neural networks, such as GCNs, that require the entire graph to be processed simultaneously, GraphSAGE introduces a more scalable approach by sampling a fixed-size neighborhood around each node and aggregating their features. This methodology allows GraphSAGE to efficiently deal with graphs of varying sizes and topologies, including those that evolve over time, by learning a function that can generate embeddings for unseen nodes based on their local neighborhoods.
The propagation rule for GraphSAGE can be generalized as:
$h^{(l+1)}_{i}=\sigma\left(W^{(l)}\cdot\text{AGGREGATE}^{(l)}\left(\{h^{(l)}_{j}|j\in\mathcal{N}(i)\}\right)\right)$ (5)
Where:
$h^{(l+1)}_{i}$ is the feature vector of node $i$ at layer $l+1$.
$\mathcal{N}(i)$ denotes the set of neighbors for node $i$.
$\text{AGGREGATE}^{(l)}$ is a function that combines the feature vectors of the sampled neighborhood nodes at layer $l$.
$W^{(l)}$ is the weight matrix for layer $l$.
$\sigma$ represents an activation function, such as the ReLU function.
GraphSAGE’s aggregation functions can vary, including mean, LSTM, and pooling aggregators, which allows the model to be tailored to specific types of graph data and applications. This flexibility, combined with the efficiency of sampling, makes GraphSAGE particularly suitable for dynamic graphs and scenarios where real-time embedding generation is crucial.
Moreover, by learning to aggregate information from a node’s local neighborhood, GraphSAGE can leverage the structural information inherent in the graph, allowing for powerful representations that capture both the features of individual nodes and their relational context within the graph. This approach has proven effective across a range of tasks, including node classification, link prediction, and graph classification, particularly in domains where the graph structure is indicative of underlying patterns or relationships, such as social networks, recommendation systems, and knowledge graphs.
6 Principal Neighborhood Aggregation
Principal Neighborhood Aggregation (PNA) addresses the need for more nuanced aggregation mechanisms capable of capturing the diverse structural properties within graphs. Diverging from traditional graph neural networks such as GCNs, which primarily utilize simplistic aggregation functions such as sum, mean, or max, PNA introduces a multifaceted approach by integrating several aggregation schemes and a degree-scaling component. This mixture of techniques enhances the model’s ability to represent the intricate patterns and relationships inherent in graph-structured data.
The core idea of PNA is to leverage the strengths of multiple aggregation functions simultaneously, thereby improving the feature representation of each node by capturing various aspects of its neighborhood’s structure and feature distribution. The inclusion of a degree-scaling component further refines this process by adjusting the influence of neighboring nodes based on their connectivity, thus providing a more balanced and informative aggregation outcome.
The general update rule for PNA can be encapsulated as:
$h^{(l+1)}_{i}=\sigma\left(\sum_{\text{agg}\in\mathcal{A}}\delta_{\text{agg}}\cdot W^{(l)}_{\text{agg}}\cdot\text{agg}\left(\{h^{(l)}_{j}|j\in\mathcal{N}(i)\cup\{i\}\}\right)\right)$ (6)
Where:
- $h^{(l+1)}_{i}$ denotes the feature vector of node $i$ at layer $l+1$.
- $\mathcal{A}$ is a set of aggregation functions, such as sum, mean, and max.
- $\delta_{\text{agg}}$ represents a degree-scaling factor associated with each aggregation function, optimizing the impact of node degrees on the aggregation process.
- $W^{(l)}_{\text{agg}}$ is a weight matrix specific to each aggregation function at layer $l$.
- $\sigma$ is an activation function, for instance, the ReLU function.
PNA’s design is particularly effective in graphs where nodes exhibit significant variability in their degree distribution. By considering multiple aggregation perspectives and adjusting for node degree, PNA ensures an accurate representation of each node’s neighborhood, significantly improving the performance on tasks such as node classification, graph classification, and link prediction.
@@ 3 Graph Neural Network Frameworks
Programming Language Primarily Python Primarily Python
Deep Learning Framework Built on top of PyTorch Supports both PyTorch and TensorFlow
Ease of Use High-level API, easy to use for beginners Low-level C++ API available
Performance Efficient and scalable Highly optimized for sparse operations
Models Available Extensive collection of pre-implemented GNN models Newer models only optimized for PyG
Community and Support Large community, active development Relatively new library with community beginning to grow
Documentation Extensive documentation with examples Comprehensive documentation
Graph Types Supported Supports heterogeneous and temporal graphs Supports heterogeneous graphs
Scalability Optimized for single-machine, multi-GPU setups Designed for distributed training on multiple machines and large graphs
Extensibility Easy to extend and contribute Low-level C++ API relatively difficult to extend
@@ 2 Prior GNN Accelerators
Accelerator Kernel Fusion Loop Reordering Pruning Bank Mapping Load Balancing
@@
ReFlip [84]: ReFlip is a GCN accelerator that aims to improve the overall efficiency of both regular neural network computations and irregular graph analytics. ReFlip employs a unified architecture based on Processing-in-Memory (PIM) [94] and features a crossbar structure. This unified architecture is augmented by novel algorithm mappings that maximize performance by leveraging the inherent parallelism of crossbar structures.
@@
LISA [115]: LISA is a compiler-oriented approach to map computations of GNNs on Coarse-Grained Reconfigurable Arrays (CGRA) spatial accelerators. CGRAs, known for their potential to enhance computational performance and energy efficiency, require sophisticated compiler designs to unlock their full capabilities. LISA introduces a solution by leveraging GNNs to analyze and interpret the structural characteristics of dataflow graphs (DFGs), which represent application-specific computations. This analysis facilitates the automatic identification of near optimal mappings for DFGs onto new accelerator architectures, considering both node placement and dependency routing. The integration of a simulated annealing-based mapping strategy, informed by GNN-generated insights, ensures that the mapping process is both efficient and effective. LISA dramatically reduces the time required to generate high-quality mappings for spatial accelerators, thereby accelerating the development cycle and enhancing the performance of computing systems.
@@
In this dissertation , we address this need by presenting GNNMark [14], a feature-rich benchmark suite that encompasses the diversity present in GNN training workloads, datasets and GNN frameworks. Our benchmark suite consists of GNN workloads that utilize various graph-based data structures, including homogeneous graphs, dynamic graphs, and heterogeneous graphs commonly used in a number of application domains that we mentioned in Section 2. We use this benchmark suite to explore and characterize GNN training behavior on GPUs. We study a variety of aspects of GNN execution, including both compute and memory behavior, highlighting major bottlenecks observed during GNN training. At the system level, we evaluate multiple metric, including the scalability of training GNNs across a multi-GPU system, as well as the sparsity of data encountered during training. The insights derived from our work can be leveraged by both hardware and software developers to improve both the hardware and software performance of GNN training on GPUs. The contributions of this part of the thesis include:
@@ 1 Motivation for Characterizing GNN Workloads
Deep Neural Networks (DNNs) have revolutionized numerous areas, such as image classification [53, 161], speech recognition [142, 49] and autonomous systems [122, 127]. Notable DNN architectures such as Convolutional Neural Networks (CNNs) [61] and Transformers [184] primarily operate on Euclidean data. This type of data, inherently 1D or 2D, includes images and speech datasets [89]. Yet, much of the data we encounter in the real world is non-Euclidean in nature [23], encompassing structures of molecules, social networks, sensor systems, and manifolds. Traditional DNNs, designed for Euclidean data, often fall short in efficiently processing non-Euclidean data due to challenges in directly applying operations, such as convolutions [23, 162].
@@
With GPUs establishing themselves as the go-to platform for DNN and GNN training due to their advanced capabilities, many leading GNN frameworks, for example, the Deep Graph Library [216] and PyTorch Geometric [58], have integrated GPU training support. As GNNs continue to surge in popularity, there’s a pressing need for optimizing GPU platforms to train them. Thoroughly analyzing GPU behavior during GNN training is paramount. By dissecting the myriad of GNN operations and their execution during training, GPU architects can pinpoint and address performance bottlenecks. Aspects such as GNN training scalability on multi-GPU setups and the presence of data sparsity during training can be tapped into for training large graphs, especially those surpassing a single GPU’s memory capacity [150, 160, 119]. A thorough analysis of GNN training workflows will enhance our understanding of the computational and memory constraints associated with running GNN workloads on GPUs.